I need to setup linear equation for the following:A 7 mile cab ride cost $10.20, a 10 mile cab ride is $13.80. Find a linear equation that models a relationship bewtween the cost (c) and the distance (d).
Answer by stanbon(75887)

I need to setup linear equation for the following:A 7 mile cab ride cost $10.20, a 10 mile cab ride is $13.80. Find a linear equation that models a relationship bewtween the cost (c) and the distance (d).
----
You have 2 points (7,10,2) and (10,13.8)
0000
slope = (13.8-10.2)/(10-7) = 3.6/3 = 1.2
---
Form: y = mx + b
You know y = 10.2 when x = 7 and m = 1.2
-------
Solve for "b":
10.2 = 1.2*7 + b
b = 1.8
----
Equation:
y = 1.2x + 1.8
===================
Cheers,
Stan H.